# Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility data

### v1

@@ Abstract
Observations with radio arrays that target the 21-cm signal originating from the early Universe suffer from a variety of systematic effects. An important class of these are reflections and spurious couplings between antennas. We apply a Hamiltonian Monte Carlo sampler to the modelling and mitigation of these systematics in simulated Hydrogen Epoch of Reionisation Array (HERA) data. This method allows us to form statistical uncertainty estimates for both our models and the recovered visibilities, which is an important ingredient in establishing robust upper limits on the Epoch of Reionisation (EoR) power spectrum. In cases where the noise is large compared to the EoR signal, this approach can constrain the systematics well enough to mitigate them down to the noise level for both systematics studied. Where the noise is smaller than the EoR, our modelling can mitigate the majority of the reflections with there being only a minor level of residual systematics, while cross-coupling sees essentially complete mitigation. Our approach performs similarly to existing filtering/fitting techniques used in the HERA pipeline, but with the added benefit of rigorously propagating uncertainties. In all cases it does not significantly attenuate the underlying signal.
@@
pubyear: 2023pagerange: Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility dataBayesian estimation of cross-coupling and reflection systematics in 21cm array visibility data
@@ 1 Introduction
Most relevant to this work, however, is the effect of internal instrument coupling. Fig. 1 shows a schematic of the systematics in HERA Phase I which are relevant to this work. As it is defined in Kern et al. (2019), these systematics are signal chain reflections, and antenna cross-coupling. Reflections result from a portion of the measured signal being reflected at either end of the coaxial cables connecting the various components in the array, for example the connections between antennas and correlators and/or digitisers. Cable wear can also result in multiple reflections as a signal travels along the cable, which is referred to as subreflections. In order to move the reflection systematic towards higher delay modes (away from those of scientific interest), HERA Phase II lengthens the connection between the Front End Module (FEM) and Post Amplifier Module (PAM) with a 500 m fiber cable (Berkhout et al. in prep.).
@@
The paper is organised as follows: Section 2 describes the systematics of interest in this work, and how they are currently removed from observational data. Section 3 details the data used, and the characteristics of the systematics added, while Section 4 discusses our foreground and EoR models, Monte Carlo sampler setup, the systematics subtraction method, and the signal loss metric. Section 5 presents the results, and Section 6 summarises the work.
@@ 2 Systematic effects
$T(\boldsymbol{\theta})$ is the sky brightness temperature, $A_{p}(\boldsymbol{\theta})$ the primary beam, $\boldsymbol{\theta}$ the coordinates on the sky in the flat-sky approximation, where $\boldsymbol{\theta}\equiv(\theta_{x},\theta_{y})$, and $\lambda$ the wavelength of the radiation measured. Most relevant is the Fourier transform (or delay transform) of this visibility:
@@ 2.1 Cable Reflections
Here, $A$ refers to the relative amplitude of the reflection in comparison to the foreground peak of the corrupted visibility, $\tau$ is its time delay, and $\phi$ is its phase offset, all for antenna 1. A reflection-corrupted cross-correlation visibility can be described as
@@
where the first term is the fiducial cross-correlation visibility formed by each antenna’s respective voltage spectrum. The second and third terms are copies of this visibility, arising from antenna 1 (at positive delays) and antenna 2 (at negative delays), respectively. The final term is a second-order effect which is weaker than the 21-cm signal, and can be safely ignored.
@@
Following the description of Aguirre et al. (2022), the total effect of $M$ cable reflections in a particular signal chain are modelled as the product of the respective gain terms,
$\tilde{g}_{a}=\prod\limits_{j}^{M}(1+A_{a,j}e^{i2\pi\nu\tau_{a,j}+i\phi_{a,j}}).$ (6)
Here, antenna $a$ is corrupted by $M$ reflections. Gain terms account for unintended variations in phase and amplitude introduced by the signal chain between the antenna and correlator (Kern et al., 2019). In this case, it can be used to correct for the excess amplitude and phase shifts produced by cable reflections.
For a given uncalibrated visibility, the reflection gains are applied with
$V_{apbq,t}^{\rm{refl}}=\tilde{g}_{a}\tilde{g}_{b}^{*}V_{apbq,t}^{\rm{uncal}},$ (7)
where $V_{apbq,t}^{\rm{uncal}}$ is the uncorrupted visibility (Aguirre et al., 2022). This refers to the visibility formed by antennas $a$ and $b$, with feeds $p$ and $q$, respectively, across an integration time of $t^{\prime}$.
@@ 2.3 Cross-coupling
The second possible source of cross-coupling is capacitive crosstalk, wherein signal chains in close proximity to one another interact electromagnetically, inducing voltages. This systematic can be mitigated at the hardware level (Chaudhari et al., 2017), and so this source of coupling is not a main concern in the current analysis, nor in Kern et al. (2019), Kern et al. (2020a), or Aguirre et al. (2022).
@@
$V_{apbq,t}^{cc}=V_{apap,t}^{\rm{refl}}\left(\sum\limits_{j}^{N}A_{apbq}^{d,j}\exp(i2\pi\nu\tau_{apbq}^{d,j}+i\phi_{apbq}^{d,j})\right)_{d\ni t},$ (10)
where the free parameters are the amplitude ($A$), delay ($\tau$), and phase ($\phi$). $V_{apbq,t}^{cc}$ is the cross-coupling spectrum which corrupts the true data, and $V_{apap,t}^{\rm{refl}}$ is the copied autocorrelation which may possibly already contain cable reflection systematics. As with the reflections, the indices $p$ and $q$ refer to the feeds of antennas $a$ and $b$, respectively, and $t^{\prime}$ the integration time. Since cross-coupling is expected to be time-stable over a night of observation, $d$ refers to a particular day. Eq. 10 only includes the cross-coupling contribution from antenna $a$, which is located at positive delays. As with all systematics analysis in this work, we only focus on those at $\tau>0$.
@@ 2.4 Current Mitigation Methods
The delay of the peak of the reflection ($\tau_{\rm peak}$) is estimated in delay space via quadratic interpolation, where an initial estimate of the peak is found by comparing the visibility amplitude at a particular delay to the amplitudes at adjacent delays. The delay corresponding to the greatest power is taken as the location of the peak. The amplitude of the reflection peak ($A$) is then estimated by the ratio of the visibility at $\tau_{\rm peak}$ and the visibility at $\tau=0$ (the foreground peak):
$A_{11}=\frac{|\tilde{V}(\tau=\tau_{\rm peak})|}{|\tilde{V}(\tau=0)|}.$ (12)
This reflection is then isolated by zeroing visibility modes on either side of this peak in order to obtain the filtered visibility $V_{\rm{filt}}$. The phase is estimated by adjusting $\phi\in[0,2\pi)$ in order to minimise the quantity $|V_{\rm{filt}}-A\exp(i2\pi\nu\tau+i\phi)|^{2}$. This provides estimates of all three parameters: $A$, $\tau$, and $\phi$. These estimates are then adjusted with a nonlinear optimisation method where they are continuously perturbed, and the reflections are iteratively calibrated out of the data using these perturbed values until the reflection peaks are minimised as much as possible.
@@
In the single cable reflection regime (where there is only one reflection peak, or peaks are clearly spaced out in delay), the Kern et al. (2019) algorithm is capable of effectively suppressing the reflection with almost no signal loss. In this single-reflection regime, approximately only 2% of the foreground and EoR signal is lost in reflection calibration. The signal loss metric is a ratio between the recovered power spectrum and the true power spectrum as a function of delay. Wherever this ratio is $<1$ is considered a loss. This metric is calculated using an ensemble average of many realisations of a sky signal, and this value of 2% is within the $1/\sqrt{N}$ sample variance of the average. In other words, this calibration does not significantly attenuate the true sky signal.
@@
These methods were applied to early HERA Phase I data (DeBoer et al., 2017) in Kern et al. (2020a). This consisted of a single 8 hour night’s worth of data, recorded with 46 antennas. Reflections in individual antennas arose from a 150 m cable connected to a post-amplifier module, followed by a 20 m cable connected to a digitiser. This data was also subject to the coupling systematics described in Section 2. A "systematic tail" was also observed, which was attributed to possible subreflections. In addition to cross-coupling, Kern et al. (2020a) were able to mitigate these systematics down to the noise floor in the power spectrum.
@@ 3 Simulated Data
This work uses the HERA Validation simulations (Aguirre et al., 2022), which include simulated point source and diffuse foregrounds, as well as simulated EoR signals. This dataset was used to validate the software pipeline used for Abdurashidova et al. (2022b). We added systematics and instrumental bandpass gains to uncorrupted datasets with hera_sim [^1] https://github.com/HERA-Team/hera_sim. These bandpass gains were not calibrated out, as we intend to follow the general approach of Kern et al. (2019) and Kern et al. (2020a) as closely as possible. The methods presented in Kern et al. (2019) make no assumption on whether instrumental bandpass has been corrected for, and Kern et al. (2020a) apply these methods to observational HERA data which contained bandpass. The authors do state, however, that bandpass calibration could lead to better defined systematics in delay space, which would in turn improve their modelling and mitigation. Our bandpass is smooth as a function of frequency, and so should not adversely affect modelling to any significant degree.
Throughout this work, then, the "true" visibilities and power spectra we intend to recover contain foregrounds, signal, and bandpass.
The Validation simulations consist of 1024 frequency channels covering 100-200 MHz. Only 512 frequency channels out of the total 1024 are used in this work, to account for any possible trimming of observational data due to RFI, for example, and to avoid requiring an excessive number of model parameters. Maintaining 512 channels should ensure that there are more data points than parameters in our model. In total, our foreground, EoR, and systematics models amount to $\sim 300-400$ parameters, but a detailed discussion is given in Section 4. Ideally, one would prefer even more observational data points in order to help constrain these parameters, but 512 is a fair compromise between modelling performance and real-world considerations.
@@
Using hera_sim, similar parameters to those in Aguirre et al. (2022) were used to generate the systematics in our data. Aguirre et al. (2022) simulated systematics matching those seen in the analysis of Kern et al. (2020a). With reference to Eq. 6, there are two high-amplitude cable reflections located at delays ($\tau$) of 200 ns and 1200 ns, with relative amplitudes ($A$) of $3\times 10^{-3}$ and $8\times 10^{-4}$, respectively. In the simulations, the delays are randomised to within 10 ns and 30 ns, respectively, and the amplitudes are randomised to within 1%. The phases ($\phi$) are randomised in the range $[-\pi,\pi)$.
Twenty cable subreflections are added, again using Eq. 6. Initially, they are evenly spaced within the delay range 200-1000 ns, but are then randomised to within 30 ns. Relative amplitudes ranged from $10^{-3}$ to $10^{-4}$ with a randomisation within 1%. Phases are again randomised within $[-\pi,\pi)$.
@@
Finally, in keeping with Eq. 10, ten cross-coupling spectra are added uniformly between delays ($\tau$) of 900-1300 ns, thereafter randomised within 20 ns. Relative amplitudes ($A$) range from $10^{-4}$ to $10^{-6}$, randomised to within 0.01% of the initial value. Phases ($\phi$) are randomised in the same manner as before. Again, we only focus on positive delays. As it was observed and mitigated in the cross-correlations of HERA Phase I data (Kern et al., 2020a), however, cross-coupling was not symmetric, with there being clear asymmetries between the positive and negative delay coupling peaks. It was determined that antenna $i$ contributed to the structure at negative delays in the cross-correlations $V_{ij}$, while antenna $j$ contributed to positive delays. The delays were a function of the distance between the leaking connection point and the antenna, and the amplitudes a function of how much a surrounding antenna’s signal was being broadcast, as well as the distance it had to travel to the antenna in question.

### v2

@@ Abstract
Observations with radio arrays that target the 21-cm signal originating from the early Universe suffer from a variety of systematic effects. An important class of these are reflections and spurious couplings between antennas. We apply a Hamiltonian Monte Carlo sampler to the modelling and mitigation of these systematics in simulated Hydrogen Epoch of Reionisation Array (HERA) data. This method allows us to form statistical uncertainty estimates for both our models and the recovered visibilities, which is an important ingredient in establishing robust upper limits on the Epoch of Reionisation (EoR) power spectrum. In cases where the noise is large compared to the EoR signal, this approach can constrain the systematics well enough to mitigate them down to the noise level for both systematics studied. Incoherently averaging the recovered power spectra can further reduce the noise and improve recovery. Where the noise level is lower than the EoR, our modelling can mitigate the majority of the reflections and coupling with there being only a minor level of residual systematics. Our approach performs similarly to existing filtering/fitting techniques used in the HERA pipeline, but with the added benefit of rigorously propagating uncertainties. In all cases it does not significantly attenuate the underlying signal.
@@
pubyear: 2023pagerange: Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility dataReferences
@@ 1 Introduction
Most relevant to this work, however, is the effect of internal instrument coupling. Fig. 1 shows a schematic of the systematics in HERA Phase I which are relevant to this work. As it is defined in Kern et al. (2019), these systematics are signal chain reflections, and antenna cross-coupling. Reflections result from a portion of the measured signal being reflected at either end of the coaxial cables connecting the various components in the array, for example the connections between antennas and correlators and/or digitisers. Cable wear can also result in multiple reflections as a signal travels along the cable, which is referred to as subreflections. In order to move the reflection systematic towards higher delay modes (away from those of scientific interest), HERA Phase II lengthens the connection between the Front End Module (FEM) and Post Amplifier Module (PAM) with a 500 m fiber cable (Berkhout et al., 2024).
@@
The paper is organised as follows: Section 2 describes the systematics of interest in this work, and how they are currently removed from observational data. Section 3 details the data used, and the characteristics of the systematics added, while Section 4 discusses our foreground model, Monte Carlo sampler setup, the systematics subtraction method, and the signal loss metric. Section 5 presents the results, and Section 6 summarises the work.
@@ 2 Systematic effects
$\rm{T}(\boldsymbol{\theta})$ is the sky brightness temperature, $\rm{A}_{\rm{p}}(\boldsymbol{\theta})$ the primary beam, $\boldsymbol{\theta}$ the coordinates on the sky in the flat-sky approximation, where $\boldsymbol{\theta}\equiv(\theta_{x},\theta_{y})$, and $\lambda$ the wavelength of the radiation measured. Most relevant is the Fourier transform (or delay transform) of this visibility:
@@ 2.1 Cable Reflections
Here, $\rm{A}$ refers to the relative amplitude of the reflection in comparison to the foreground peak of the corrupted visibility, $\tau$ is its time delay, and $\phi$ is its phase offset, all for antenna 1. A reflection-corrupted cross-correlation visibility can be described as
@@
where the first term is the fiducial cross-correlation visibility formed by each antenna’s respective voltage spectrum. The second and third terms are copies of this visibility, arising from antenna 1 (at positive delays) and antenna 2 (at negative delays), respectively. The final term is a second-order effect. Depending on the strength of the coupling coefficient, this may produce an additional non-negligible systematic feature at higher delays, or it could result in a systematic which is weaker than the 21-cm signal.
For the systematics simulated in this work, first-order reflections begin at $200$ ns. With relative amplitudes of $\sim 10^{-3}$ (in visibility-space), second-order reflections would begin at $400$ ns, and would be slightly stronger than the foregrounds and signal. Similar can be said of the power of the second-order coupling systematics. It would be fairly straightforward to include second-order reflections, as the delays are simply dependent on both antenna one and antenna two’s cable lengths, and the relative amplitudes would be the product of $A_{11}$ and $A_{22}$. However, noise and imperfect knowledge of the instrument would complicate matters, so distinct parameters for first and second-order effects would possibly be required. It is certainly feasible, though, to use the model estimates for the first-order systematics to estimate the second-order parameters, rather than sampling for the latter. At the very least the relation between the different orders can be used to inform the priors. We opt to neglect the second-order terms, however, in order to avoid additional parameters in an already high-dimensional model.
@@
Following the description of Aguirre et al. (2022), the total effect of $\rm{M}$ cable reflections in a particular signal chain are modelled as the product of the respective gain terms,
$\tilde{g}_{a}=\prod\limits_{j}^{\rm{M}}(1+{\rm{A}}_{a,j}e^{i2\pi\nu\tau_{a,j}+i\phi_{a,j}}).$ (6)
Here, antenna $a$ is corrupted by $\rm{M}$ reflections. Gain terms account for unintended variations in phase and amplitude introduced by the signal chain between the antenna and correlator (Kern et al., 2019). In this case, it can be used to correct for the excess amplitude and phase shifts produced by cable reflections.
For a given uncorrupted visibility, the reflection gains are applied with
$V_{apbq,t}^{\rm{refl}}=\tilde{g}_{a}\tilde{g}_{b}^{*}V_{apbq,t}\>,$ (7)
where $V_{apbq,t}$ is the true visibility (Aguirre et al., 2022). This refers to the visibility formed by antennas $a$ and $b$, with feeds $p$ and $q$, respectively, across an integration time of $t$.
@@ 2.3 Cross-coupling
The second possible source of cross-coupling is capacitive crosstalk, wherein signal chains in close proximity to one another interact electromagnetically, inducing voltages. This systematic can be mitigated at the hardware level (Chaudhari et al., 2017), and so this source of coupling is not a concern in the current analysis, nor in Kern et al. (2019), Kern et al. (2020a), or Aguirre et al. (2022).
@@
$V_{apbq,t}^{cc}=V_{apap,t}^{\rm{refl}}\left(\sum\limits_{j}^{\rm{N}}{\rm{A}}_{apbq}^{d,j}\exp(i2\pi\nu\tau_{apbq}^{d,j}+i\phi_{apbq}^{d,j})\right)_{d\ni t},$ (10)
where $\rm{N}$ is the number of cross-coupling peaks present in the visibility, and the free parameters are the amplitude ($\rm{A}$), delay ($\tau$), and phase ($\phi$). $V_{apbq,t}^{cc}$ is the cross-coupling spectrum which corrupts the true data, and $V_{apap,t}^{\rm{refl}}$ is the copied autocorrelation which may possibly already contain cable reflection systematics. As with the reflections, the indices $p$ and $q$ refer to the feeds of antennas $a$ and $b$, respectively, and $t$ the integration time. Since cross-coupling is expected to be time-stable over a night of observation, $d$ refers to a particular day. Eq. 10 only includes the cross-coupling contribution from antenna $a$, which is located at positive delays. As with all systematics analysis in this work, we only focus on those at $\tau>0$.
@@ 2.4 Current Mitigation Methods
The delay of the peak of the reflection ($\tau_{\rm peak}$) is estimated in delay space via quadratic interpolation, where an initial estimate of the peak is found by comparing the visibility amplitude at a particular delay to the amplitudes at adjacent delays. The delay corresponding to the greatest power is taken as the location of the peak. The amplitude of the reflection peak ($\rm{A}$) is then estimated by the ratio of the visibility at $\tau_{\rm peak}$ and the visibility at $\tau=0$ (the foreground peak):
$\rm{A}_{11}=\frac{|\tilde{V}(\tau=\tau_{\rm peak})|}{|\tilde{V}(\tau=0)|}.$ (12)
This reflection is then isolated by zeroing visibility modes on either side of this peak in order to obtain the filtered visibility $V_{\rm{filt}}$. The phase is estimated by adjusting $\phi\in[0,2\pi)$ in order to minimise the quantity $|V_{\rm{filt}}-\rm{A}\exp(i2\pi\nu\tau+i\phi)|^{2}$. This provides estimates of all three parameters: $\rm{A}$, $\tau$, and $\phi$. These estimates are then adjusted with a nonlinear optimisation method where they are continuously perturbed, and the reflections are iteratively calibrated out of the data using these perturbed values until the reflection peaks are minimised as much as possible.
@@
To improve the resistance to signal loss, the $\mathbf{T}$ modes are low-pass filtered, with a Gaussian process regression filter providing the best results. This ensures that only low-fringe-rate modes (where the coupling systematic is predominantly located) undergo subtraction, while the signal itself dominates high fringe rates.
This could not be immediately replicated with our method as we do not form similar modes in our model. We instead form model visibilities which are compared to the mock data. The closest approximation would be, for multiple integrations, to filter our model cross-coupling-only visibilities across time, and using these to subtract this systematic from the mock data. This would possibly eliminate high-fringe-rate structure in our coupling model, as it does in Kern et al. (2019)’s method. However, our intent is to assess the viability of forward-modelling the systematics on a per-antenna, per-time-integration basis, without any additional filtering, SVD approaches, etc.
In the single cable reflection regime (where there is only one reflection peak, or peaks are clearly spaced out in delay), the algorithm of Kern et al. (2019) is capable of effectively suppressing the reflection with almost no signal loss. In this single-reflection regime, approximately 2% of the foreground and EoR signal is lost in reflection calibration. The signal loss metric is a ratio between the recovered power spectrum and the true power spectrum as a function of delay. Wherever this ratio is $<1$ is considered a loss. This metric is calculated using an ensemble average of many realisations of a sky signal, and this value of 2% is within the ${\rm{N}}^{-\frac{1}{2}}$ sample variance of the average. In other words, this calibration does not significantly attenuate the true sky signal.
@@
These methods were applied to early HERA Phase I data (DeBoer et al., 2017) in Kern et al. (2020a). This consisted of a single 8 hour night’s worth of data, recorded with 46 antennas. Reflections in individual antennas arose from a 150 m cable connected to a post-amplifier module, followed by a 20 m cable connected to a digitiser. This data was also subject to the coupling systematics described in Section 2. A “systematic tail” was also observed, which was attributed to possible subreflections. In addition to cross-coupling, Kern et al. (2020a) were able to mitigate these systematics down to the noise floor in the power spectrum.
@@ 3 Simulated Data
This work uses the HERA Validation simulations (Aguirre et al., 2022), which include simulated point source and diffuse foregrounds, as well as simulated EoR signals. This dataset was used to validate the software pipeline used in Abdurashidova et al. (2022b). We added systematics and instrumental bandpass gains to uncorrupted datasets with hera_sim [^1] https://github.com/HERA-Team/hera_sim. The bandpass gains are randomly generated on a per-antenna basis, and as a function of frequency. Starting with the default HERA H1C bandpass in hera_sim, randomised white noise multiplied by a spread factor (10% being the default) is then added. This spread factor is the standard deviation of the randomised gains from antenna to antenna, i.e. across all antennas in the array, the generated bandpass gains vary by up to a factor of 0.1.
These bandpass gains were not calibrated out, as we intend to follow the general approach of Kern et al. (2019) and Kern et al. (2020a) as closely as possible. The methods presented in Kern et al. (2019) make no assumption on whether instrumental bandpass has been corrected for, and Kern et al. (2020a) apply these methods to observational HERA data which contained bandpass. The authors do state, however, that bandpass calibration could lead to better defined systematics in delay space, which would in turn improve their modelling and mitigation. Our bandpass is smooth as a function of frequency, and so should not adversely affect modelling to any significant degree.
Throughout this work, then, the “true” visibilities and power spectra we intend to recover contain foregrounds, signal, and bandpass.
The Validation simulations consist of 1024 frequency channels covering 100–200 MHz. Only 512 frequency channels out of the total 1024 are used in this work, to account for any possible trimming of observational data due to RFI, for example, and to avoid requiring an excessive number of model parameters. Maintaining 512 channels should ensure that there are more data points than parameters in our model. In total, our foreground and systematics models amount to $\sim 100$ parameters. A detailed discussion is given in Section 4. Ideally, one would prefer even more observational data points in order to help constrain these parameters, but 512 is a fair compromise between modelling performance and real-world considerations.
@@
Using hera_sim, similar parameters to those in Aguirre et al. (2022) were used to generate the systematics in our data. Aguirre et al. (2022) simulated systematics matching those seen in the analysis of Kern et al. (2020a).
With reference to Eq. 6, there are two high-amplitude cable reflections located at delays ($\tau$) of 200 ns and 1200 ns, with relative amplitudes ($\rm{A}$) of $3\times 10^{-3}$ and $8\times 10^{-4}$, respectively. In the simulations, the delays are randomised to within 10 ns and 30 ns, respectively, and the amplitudes are randomised to within 1%. The phases ($\phi$) are randomised in the range $[-\pi,\pi)$.
Twenty cable subreflections are added, again using Eq. 6. Initially, they are evenly spaced within the delay range 200–1000 ns, but are then randomised to within 30 ns. Relative amplitudes ranged from $10^{-3}$ to $10^{-4}$ with a randomisation within 1%. Phases are again randomised within $[-\pi,\pi)$.
@@
Finally, in keeping with Eq. 10, ten cross-coupling spectra are added uniformly between delays ($\tau$) of 900-1300 ns, thereafter randomised within 20 ns. Relative amplitudes ($\rm{A}$) range from $10^{-4}$ to $10^{-6}$, randomised to within 0.01% of the initial value. Phases ($\phi$) are randomised in the same manner as before. Again, we only focus on positive delays. As it was observed and mitigated in the cross-correlations of HERA Phase I data (Kern et al., 2020a), however, cross-coupling was not symmetric, with there being clear asymmetries between the positive and negative delay coupling peaks. It was determined that antenna $i$ contributed to the structure at negative delays in the cross-correlations $V_{ij}$, while antenna $j$ contributed to positive delays. The delays were a function of the distance between the leaking connection point and the antenna, and the amplitudes a function of how much a surrounding antenna’s signal was being broadcast, as well as the distance it had to travel to the antenna in question.